OpenStudy (anonymous):

Find the derivative of cos^2(3sqrt(x)) For my answer I get (6cos*3sqrt(x))/2sqrt(x) But the book says the answer is (-3/sqrt(x))*cos(3sqrtx)*sin(3sqrtx)

OpenStudy (phi):

\[ \frac{d}{dx} u^2 = 2 u \frac{d}{dx}u \] if u is cos(stuff) \[ \frac{d}{dx} \cos(stuff)^2 = 2 \cos(stuff) \frac{d}{dx}\cos(stuff) \] we know the derivative of cos is sin (the rule is): \[ \frac{d}{dx} \cos(u) = \sin(u) \frac{d}{dx}u \] so you definitely should get a sin in your answer.